### v3

@@ 3.1 Notation
Now, let $p=1,...,P$ index the practices and let $\mathds{1}_{{p,i}}$ stand for the practice assignment indicator of patient $i$. Write $Y_{i}(p)$ for the potential outcome (Neyman 1923, 1990; Rubin 1974) under practice assignment $\mathds{1}_{{p,i}}$, $p=1,...,P$. Denote $Y_{i}$ for their observed outcome, $Y_{i}=\sum_{p=1}^{P}\mathds{1}_{{p,i}}Y_{i}(p)$. Finally, define $\boldsymbol{U}$ as the vector unobserved covariates. See Table A1 in the Online Supplementary Materials for a summary of this notation.
@@ 4 Methods for layered case-mix adjustments
The methods we consider are summarized in Table A2 in the Online Supplementary Materials. They combine weighting and modeling approaches or use one of them. First, we describe traditional modeling approaches. Afterward, we consider approaches that start by weighting and increasingly impose stronger parametric modeling assumptions where there is less or no data.
@@
For weighting-only approaches, we consider inverse probability weighting using logistic regression (LW) and the stable balancing weighting (SBW; Zubizarreta 2015). Unlike modeling approaches to weighting, the latter balancing approach is better suited for comparing hundreds of treatments (i.e., practices in our case study). Specifically, SBW finds weights of minimum variance that approximately balance functions of the covariates. See Wang and Zubizarreta (2020) and Chattopadhyay et al. (2020) for theoretical properties and empirical performance of the SBW in causal inference settings. Ideally, the weights will take non-negative values and balance the covariates by interpolating within the convex hull of the available data. Of course, this non-extrapolating form of adjustment is feasible only for practices that contain $\boldsymbol{x}^{*}$ inside its convex hull; otherwise, the optimization problem does not admit a solution indicating that the investigator needs to extrapolate to adjust. In Table A2 (see the Online Supplementary Materials), we denote this approach by $\{\textrm{SBW}(\mathbb{R}^{+}_{0},\boldsymbol{X}),\cdot\}$ or $\{\textrm{SBW}(\mathbb{R}^{+}_{0},\widetilde{\boldsymbol{X}}),\cdot\}$ depending on whether the original covariates $\boldsymbol{X}$ or their transformations $\widetilde{\boldsymbol{X}}$ are used (in our empirical studies, $\widetilde{\boldsymbol{X}}$ includes the square of principal component scores of the covariates in order to approximately balance their covariance matrix). Since some covariates are null for some of the practices (that is, when some covariates take a constant value different to the corresponding values of $\boldsymbol{x}^{*}$ for all patients in some of the practices), alternatively, we combine the SBW with a modeling step as follows.
@@ 5.2 Comparative performances
We evaluate the performance of all the methods from two perspectives: an absolute perspective, focusing on bias and root mean square error (RMSE) of the estimated survival rates (Tables 1 and 2), and a relative perspective, based on the errors in the estimated practice rankings (Tables A3 and A4). For both perspectives, we consider three target profiles: patients across the overall system of practices, patients in one practice near the center of covariate distribution, and patients in one practice at the boundary of covariate distribution. We evaluate the performance of the methods described in Table A2.
@@
In Tables A3 and A4, two metrics are presented: the mean and the maximum absolute difference between the true and estimated rankings. For both metrics, the findings are consistent with the results observed previously.
@@ 6 Case study
For each of these profiles, we inspect the data to understand the extent to which data are available to assess the survival outcomes for each of the different practices. Figure A1 in the Online Supplementary Materials counts the number of practices with null-case covariates for each profile. Again, these are practices for which it is not possible to carry out the case-mix adjustments (and therefore, estimate the mean outcome) using data solely from the practice in question; that is, without leveraging the functional form of a model fitted on data from patients in other practices. For the system profile, across all 600 practices, there are 306 practices with no null-case covariates and 294 practices with at least one null-case covariate. Among these, 83 have three or more null-case covariates. The covariates that produce the largest number of null cases are race (Black and other), education (quartile 1), income (quartile 4), and cancer type (ovary and prostate). For patients 1, 2, and 3, the numbers of practices with null-case covariates are reduced substantially, but there are still sizable numbers of practices that did not treat a single patient with a certain covariate value of the target profiles. This is particularly true for Patient 3, because many practices did not treat any Black women with pancreatic cancer, a cancer that is relatively infrequently diagnosed.
@@
Figure A2 in the Online Supplementary Materials shows the changes in rankings for different profiles. The left plot compares the rankings for the system profile and that of Patient 1. There are substantial differences: 367 or 61.1% of the practices change their ranking by 10% (60 ranking places) or more. The center plot makes a similar comparison between profiles 2 and 3, when only one covariate value (race) changes between profiles. 426 or 71% of the practices change their ranking by 10% (60 ranking places) or more. To better grasp this variation, the right plot shows the changes in relative positions after grouping the practices by quintiles. Table 4 counts the number of practices for each of these changes: 176 or 29.3% of the practices remain in the same quintile when the profile changes (diagonal of the matrix), 215 or 35.8% of the practices move up or down by one quintile, 209 or 34.8% move by two or more quintiles, and 42 or 7% of the practices move from the top to the bottom quintile or vice versa. This exemplifies that the profile — the “test” on which each practice is evaluated — matters. Among other reasons, this is important for quality measurement that is used for high-stakes purposes, such as public reporting or payment.

### v4

@@ 3.1 Notation
Now, let $p=1,...,P$ index the practices and let $\mathds{1}_{{p,i}}$ stand for the practice assignment indicator of patient $i$. Write $Y_{i}(p)$ for the potential outcome (Neyman 1923, 1990; Rubin 1974) under practice assignment $\mathds{1}_{{p,i}}$, $p=1,...,P$. Denote $Y_{i}$ for their observed outcome, $Y_{i}=\sum_{p=1}^{P}\mathds{1}_{{p,i}}Y_{i}(p)$. Finally, define $\boldsymbol{U}$ as the vector unobserved covariates. See Table LABEL:tab_notation in the Online Supplementary Materials for a summary of this notation.
@@ 4 Methods for layered case-mix adjustments
The methods we consider are summarized in Table LABEL:tab_methods in the Online Supplementary Materials. They combine weighting and modeling approaches or use one of them. First, we describe traditional modeling approaches. Afterward, we consider approaches that start by weighting and increasingly impose stronger parametric modeling assumptions where there is less or no data.
@@
For weighting-only approaches, we consider inverse probability weighting using logistic regression (LW) and the stable balancing weighting (SBW; Zubizarreta 2015). Unlike modeling approaches to weighting, the latter balancing approach is better suited for comparing hundreds of treatments (i.e., practices in our case study). Specifically, SBW finds weights of minimum variance that approximately balance functions of the covariates. See Wang and Zubizarreta (2020) and Chattopadhyay et al. (2020) for theoretical properties and empirical performance of the SBW in causal inference settings. Ideally, the weights will take non-negative values and balance the covariates by interpolating within the convex hull of the available data. Of course, this non-extrapolating form of adjustment is feasible only for practices that contain $\boldsymbol{x}^{*}$ inside its convex hull; otherwise, the optimization problem does not admit a solution indicating that the investigator needs to extrapolate to adjust. In Table LABEL:tab_methods (see the Online Supplementary Materials), we denote this approach by $\{\textrm{SBW}(\mathbb{R}^{+}_{0},\boldsymbol{X}),\cdot\}$ or $\{\textrm{SBW}(\mathbb{R}^{+}_{0},\widetilde{\boldsymbol{X}}),\cdot\}$ depending on whether the original covariates $\boldsymbol{X}$ or their transformations $\widetilde{\boldsymbol{X}}$ are used (in our empirical studies, $\widetilde{\boldsymbol{X}}$ includes the square of principal component scores of the covariates in order to approximately balance their covariance matrix). Since some covariates are null for some of the practices (that is, when some covariates take a constant value different to the corresponding values of $\boldsymbol{x}^{*}$ for all patients in some of the practices), alternatively, we combine the SBW with a modeling step as follows.
@@ 5.2 Comparative performances
We evaluate the performance of all the methods from two perspectives: an absolute perspective, focusing on bias and root mean square error (RMSE) of the estimated survival rates (Tables 1 and 2), and a relative perspective, based on the errors in the estimated practice rankings (Tables LABEL:tab_the_practice_perspective and LABEL:tab_the_practice_perspective_200). For both perspectives, we consider three target profiles: patients across the overall system of practices, patients in one practice near the center of covariate distribution, and patients in one practice at the boundary of covariate distribution. We evaluate the performance of the methods described in Table LABEL:tab_methods.
@@
In Tables LABEL:tab_the_practice_perspective and LABEL:tab_the_practice_perspective_200, two metrics are presented: the mean and the maximum absolute difference between the true and estimated rankings. For both metrics, the findings are consistent with the results observed previously.
@@ 6 Case study
For each of these profiles, we inspect the data to understand the extent to which data are available to assess the survival outcomes for each of the different practices. Figure LABEL:fig_null_covariates in the Online Supplementary Materials counts the number of practices with null-case covariates for each profile. Again, these are practices for which it is not possible to carry out the case-mix adjustments (and therefore, estimate the mean outcome) using data solely from the practice in question; that is, without leveraging the functional form of a model fitted on data from patients in other practices. For the system profile, across all 600 practices, there are 306 practices with no null-case covariates and 294 practices with at least one null-case covariate. Among these, 83 have three or more null-case covariates. The covariates that produce the largest number of null cases are race (Black and other), education (quartile 1), income (quartile 4), and cancer type (ovary and prostate). For patients 1, 2, and 3, the numbers of practices with null-case covariates are reduced substantially, but there are still sizable numbers of practices that did not treat a single patient with a certain covariate value of the target profiles. This is particularly true for Patient 3, because many practices did not treat any Black women with pancreatic cancer, a cancer that is relatively infrequently diagnosed.
@@
Figure LABEL:fig_rankings in the Online Supplementary Materials shows the changes in rankings for different profiles. The left plot compares the rankings for the system profile and that of Patient 1. There are substantial differences: 367 or 61.1% of the practices change their ranking by 10% (60 ranking places) or more. The center plot makes a similar comparison between profiles 2 and 3, when only one covariate value (race) changes between profiles. 426 or 71% of the practices change their ranking by 10% (60 ranking places) or more. To better grasp this variation, the right plot shows the changes in relative positions after grouping the practices by quintiles. Table 4 counts the number of practices for each of these changes: 176 or 29.3% of the practices remain in the same quintile when the profile changes (diagonal of the matrix), 215 or 35.8% of the practices move up or down by one quintile, 209 or 34.8% move by two or more quintiles, and 42 or 7% of the practices move from the top to the bottom quintile or vice versa. This exemplifies that the profile — the “test” on which each practice is evaluated — matters. Among other reasons, this is important for quality measurement that is used for high-stakes purposes, such as public reporting or payment.